A parallelogram is a quadrilateral with two pairs of parallel sides. Show that ABCD is a parallelogram if the coordinates of the vertices are A(-2,3), B(2,7), C(8,5), and D(4,1).
Answer by checkley71(8403)

TO PROVE THIS IS A PARALLELOGRAM WE NEED TO PROVE THAT THE SLOPESOF THE OPOSITE SIDES ARE EQUAL.
(7-3)/2+2)=4/4=1 & (1-5)/(4-8)=-4/-4=1 THESE 2 LINES ARE PARALLEL.
(1-3)/(4+2)=-2/6=-3 & (5-7)/(8-2)=-2/6=-3 THESE 2 LINES ARE PARALLEL.